XYZ Co. operates in a competitive market. Its Total Product (Q) is given as Q=f(L)= 3L, and it takes the wage and price as given. Derive the firm's short-run demand for labor as a function of w and p. How much labor will the firm hire if W= ¢25 and P=¢ 150?

1
Expert's answer
2021-11-03T11:39:19-0400


When you divide the nominal demand by the product price p"\times" MRPL=w you will find the short-run labor demand curve.

It is given by MPL=W/p=w

MPL "\times" P = w


MPL = 3L

P = 150

W = 25

3L "\times" 150 = 25

450L = 25

L = 0.056

Q = 3L = 3 "\times" 0.056 = 0.168

The firm will hire labor=0.168